Cash flow of accounts receivable. Myers and​ Associates, a famous law firm in​ California, bills its clients on the first of each month. Clients pay in the following​ fashion: ​ 40% pay at the end of the first​ month, 30% pay at the end of the second​ month, 20% pay at the end of the third​ month, 5% pay at the end of the fourth​ month, and​ 5% default on their bills. Myers wants to know the anticipated cash flow for the first quarter of 2015 if the past billings and anticipated billings follow this same pattern. The actual and anticipated billings are as​ follows:

Fourth Quarter Actual Billings

First Quarter Anticipated Billings

Oct.

Nov.

Dec.

Jan.

Feb.

Mar.

​$290,000

​$236,000

​$214,000

​$246,000

​$267,000

​$298,000

What is the anticipated cash flow for January of 2015 if past billings and anticipated billings follow this same​ pattern? $________ ​(Round to the nearest​ dollar.)

What is the anticipated cash flow for February of 2015 if past billings and anticipated billings follow this same​ pattern? $_______ (Round to the nearest​ dollar.)

What is the anticipated cash flow for March of 2015 if past billings and anticipated billings follow this same​ pattern? ​$_______ ​(Round to the nearest​ dollar.)
Answer #1

Anticipated Cash flow for Month of january:

40% of billing of this month(january)

(246000*40%)= 98400

30% of previous month(December)

(214000*30%)=. 64200

20% of previous second month (November)

(236000*20%)= 47200

5% of Previous third month

(290000*5%)= 14500

____'__________________________________________

224300

________________________________________________

Anticipated Cash flow for Month of february:

40% of billing of this month(February)

(267000*40%)= 106800

30% of previous month(January)

(246000*30%)= 73800

20% of previous second month (December)

(214000*20%)= 42800

5% of Previous third month(November)

(236000*5%)=. 11800

__________________________________

235200

__________________________________

Anticipated Cash flow for Month of March:

40% of billing of this month(March)

(298000*40%)=. 119200

30% of previous month(February)

(267000*30%)= 80100

20% of previous second month (january)

(246000*20%)=. 49200

5% of Previous third month(November)

(214000*5%)= 10700

__________________________________

259200

__________________________________

So, Anticipated Cash flow for Month of january is $ 224300

Anticipated Cash flow for Month of february is $235200

Anticipated Cash flow for Month of March is $259200

Please thumbs up.